OpenStudy (anonymous):

find a radical equation of the form sqrt(ax+b)=x+c so that one solution is extraneous

OpenStudy (anonymous):

okay. give an example of square root of a number.

OpenStudy (anonymous):

4

OpenStudy (anonymous):

okay so 4 is the square root of 16. so ax+b =16 and x+c = 4.

OpenStudy (anonymous):

say x is 3, then c is 1. and 3a+b is 16. lets say a is 4, then b is also 4.

OpenStudy (anonymous):

so your equation could be sqrt(4x+4)=x+1

OpenStudy (anonymous):

to find x, square both sides. and post what you got.

OpenStudy (anonymous):

-x^2+4x+3=0

OpenStudy (anonymous):

x=-1,-3

OpenStudy (anonymous):

wait. what is (x+1)^2?

OpenStudy (anonymous):

x^2+1

OpenStudy (anonymous):

(a+b)^2 is a^2+b^2? are you sure about that?

OpenStudy (anonymous):

its 4x+4=x^2+1^2

OpenStudy (anonymous):

do you know the formula for \[(a+b)^{2} \]

OpenStudy (anonymous):

okay, if you don't know the formula, you can tell me. I will explain.

OpenStudy (anonymous):

wait i no this its 4x+4=x^2+2x+1

OpenStudy (anonymous):

yes! solve that expression then.

OpenStudy (anonymous):

(x=-1,3)

OpenStudy (anonymous):

y, plug in x = 3 and x = -1 in your original expression, sqrt(4x+4)=x+1

OpenStudy (anonymous):

oh wait. this is a bad example.

OpenStudy (anonymous):

well, anyway, you get the idea. you have to follow the procedure I outlines, such that the original expression is solve by one value of x and the square expression is solved by 2 values of x, giving you one spurious value of x.